OpenStudy (anonymous):

1/49 + 3/14 A) determine the LCD using prime factorization B) convert the original fractions to the equivilant expressions by converting the denominators to the LCD C) add the equvilant expressions and reduce if possible

OpenStudy (anonymous):

49 = 7 * 7 14 = 2 * 7 lcd = 49 * 2 = 98

OpenStudy (mathstudent55):

Prime factorization is finding the prime factors of each number. 49 = 7^2 14 = 2 * 7 To get the LCD, you need common and not common factors with the highest power. 7 is a common factor, so pick 7^2 (higher power than just 7) 2 is a not-common factor, so pick it too. LCD = 7^2 * 2 = 49 * 2 = 98

OpenStudy (mathstudent55):



OpenStudy (anonymous):

would b) be 1/98 +3/98 ?

OpenStudy (mathstudent55):

No. Did you notice the two questions at the bottom of the drawing?